A 3.2 kg sloth hangs 3.2 m above the ground. (a) What is the gravitational potential energy of the sloth-Earth system if we take the reference point y = 0 to be at the ground? J (b) Assume the sloth drops to the ground and air drag on it is negligible. What is the kinetic energy of the sloth just before it reaches the ground? J (c) What is the speed of the sloth at this time?
Answer #1

a) gravitational potential energy=mgh=3.2*2.3*9.8=100.352 J

b) kinetic energy of the sloth just before it reaches the ground=[1/2]mv^2=[1/2]m*2hg=mhg=100.352 J

c) speed of sloth=srqt(2hg)=7.92 m/s